1. y'' + 4y' + 7y = 0      ans, y= e-2x(c1cosx+c2sinx)
  2. 2y'' - 7y' + 3y = 0     ans, y=c1e3x+c2e1/2x
1
Expert's answer
2021-12-07T13:46:43-0500

1.

Corresponding (auxiliary) equation


"r^2+4r+7=0"

"D=(4)^2-4(1)(7)=-12"

"r=\dfrac{-4\pm\sqrt{-12}}{2(1)}=-2\pm i\sqrt{3}"

The general solution of the given differential equation is


"y=e^{-2x}(c_1\cos(\sqrt{3}x)+c_2\sin(\sqrt{3}x))"

2.

Corresponding (auxiliary) equation


"2r^2-7r+3=0"

"D=(-7)^2-4(2)(3)=25"

"r=\dfrac{7\pm\sqrt{25}}{2(2)}=\dfrac{7\pm5}{4}"

"r_1=\dfrac{7+5}{4}=3, r_2=\dfrac{7-5}{4}=\dfrac{1}{2}"

The general solution of the given differential equation is


"y=c_1e^{3x}+c_2e^{x\/2}"